### v2

# CodeTaxo: Enhancing Taxonomy Expansion with Limited Examples via Code Language Prompts

Abstract
Taxonomies provide structural representations of knowledge and are crucial in various applications. The task of taxonomy expansion involves integrating emerging entities into existing taxonomies by identifying appropriate parent entities for these new query entities. Previous methods rely on self-supervised techniques that generate annotation data from existing taxonomies but are less effective with small taxonomies (fewer than 100 entities). In this work, we introduce CodeTaxo, a novel approach that leverages large language models through code language prompts to capture the taxonomic structure. Extensive experiments on five real-world benchmarks from different domains demonstrate that CodeTaxo consistently achieves superior performance across all evaluation metrics, significantly outperforming previous state-of-the-art methods. The code and data are available at https://github.com/QingkaiZeng/CodeTaxo-official.
1 Introduction
Taxonomies are hierarchical structures encoding hypernym–hyponym (i.e., “is-A”) relations between concepts or entities. Relational knowledge derived from taxonomies has been widely leveraged to identify semantic relevance for web search Yin and Shah (2010); Liu et al. (2020); Kang et al. (2024), personalized recommendation Zhang et al. (2014); Tan et al. (2022); Huang et al. (2019), and question answering Yang et al. (2017). However, existing taxonomies are mainly constructed by experts or through crowd-sourcing, making the process time-consuming, labor-intensive, and restricted in coverage Bordea et al. (2016); Jurgens and Pilehvar (2016). As new entities emerge, continually enriching taxonomies with these additions becomes vital. To address these challenges, taxonomy expansion aims to integrate new entities into existing taxonomies automatically.
(a) Discriminative Methods: A trained scoring function selects the most appropriate parent entity from the taxonomy for a given query entity.
(b) Generative Methods: LLMs generate the parent entity from the taxonomy based on the query entity.
Figure 1: Two Types Taxonomy Expansion Methods
Figure 2: The overview of the pipeline for CodeTaxo: CodeTaxo reformulates the task of integrating a query entity $q$ into an existing taxonomy $\mathcal{T}_{0}$ as a code completion task using code-based prompts for LLMs.
As shown in Figure 1(a), recent taxonomy expansion methods mainly rely on discriminative methods that model hierarchical structures through techniques like Egonets Shen et al. (2020), mini-paths Yu et al. (2020), and Ego-Trees Wang et al. (2021). Although pre-trained language models (PLMs) enhance these methods by encoding entities’ textual descriptions Wang et al. (2021, 2022); Liu et al. (2021b); Xu et al. (2022), their reliance on limited self-supervised annotations often restricts performance. In contrast, Generative Large Language Models (LLMs) such as GPT-4 Achiam et al. (2023) and Llama family Touvron et al. (2023); Dubey et al. (2024) have recently shown remarkable capabilities in text comprehension and generation, making them highly effective for tasks aimed at generating structural knowledge Ye et al. (2022); Bi et al. (2024); Sun et al. (2024a, b). Increasing LLM parameters boosts generalization, surpassing smaller models and enabling superior few-shot or zero-shot performance. Even with limited annotations, LLMs effectively leverage extensive knowledge embedded within their parameters, acquired from large-scale pre-training corpora. In Figure 1(b), we illustrate the pipeline to how generative methods are applied to the taxonomy expansion task.
Two key challenges arise when applying LLMs to taxonomy expansion. First, unlike traditional text-to-text NLP tasks such as question answering and machine translation, representing the taxonomic structure for this task in natural language is inherently challenging. Specifically, the process requires "flattening" the taxonomy into a sequence of parent-child entity pairs Madaan et al. (2022), effectively serializing a hierarchical structure into linear text. This serialized format is notably different from the unstructured text that LLMs primarily encounter during pre-training. Furthermore, while semantically related words in natural language are usually located near each other, linearizing a taxonomy can separate conceptually related entities by significant distances within the sequence. This disparity adds to the difficulty of aligning LLM outputs with the desired structured representation. Second, scaling to large taxonomies amplifies the problem, as including every entity from the existing taxonomy in the prompt is infeasible. The limited contextual window size of current LLMs and the associated computational overhead imposes strict constraints. Even if it were possible to include thousands of entities within a prompt, the resulting structural information loss would impair the clarity of entity-specific distinctions, reducing the model’s capacity to effectively utilize the taxonomy.
To overcome these challenges, we propose CodeTaxo, a novel taxonomy expansion approach that leverages code language as prompts. Code-based representations have shown promise in structure prediction tasks Madaan et al. (2022); Li et al. (2023); Wang et al. (2023); Li et al. (2024); Bi et al. (2024), as code languages provide a more natural format for structural data. In CodeTaxo, we frame taxonomy expansion as a code completion task. We introduce a base Entity class to store entity surface names, definitions, parent references, and child lists, along with two methods for modifying the taxonomic relations between entities. Each existing taxonomy entity is instantiated as a corresponding Entity object. Due to constraints of contextual window size, we apply a similarity-based filter, using SimCSE Gao et al. (2021) to encode textual description for entities, to include only the most relevant entities in the prompt
We evaluate CodeTaxo through extensive experiments on two sets of small-scale WordNet and Graphine sub-taxonomies Bansal et al. (2014); Liu et al. (2021a), as well as three large-scale SemEval-2016 taxonomies Bordea et al. (2016). Our one-shot CodeTaxo surpasses all self-supervised baselines trained on large-scale SemEval-2016 annotations, achieving relative accuracy improvements of 10.26%, 8.89%, and 9.21% on SemEval-Sci, SemEval-Env, and SemEval-Food, respectively. Additionally, we evaluated CodeTaxo using various open-source LLMs, revealing several interesting observations discussed in this work.
In summary, our main contributions include:
  • We introduce CodeTaxo, an innovative in-context learning method that utilizes code language prompts to represent taxonomic relationships between entities, thereby improving the effectiveness of taxonomy expansion.
  • We develop a similarity-based filter, which employs a small pre-trained model to encode the textual descriptions of entities, ensuring that only highly relevant entities are included in the prompt concerning the query entity.
  • Extensive experiments demonstrate that CodeTaxo significantly enhances the performance of taxonomy expansion across two sets of small-scale sub-taxonomies and three large-scale taxonomies.
2 Problem Definition
Figure 3: Prompt Overview of CodeTaxo
Definition 1 (Taxonomy)
We follow the definition of taxonomy in Jiang et al. (2023). A taxonomy $\mathcal{T}=(\mathcal{E},\mathcal{H})$ is a tree-like structure, where each entity $e\in\mathcal{E}$ is a conceptual entity, and each edge $h\in\mathcal{H}$ represents the hypernymy-hyponymy relation between the two entities connected by it. Each entity $e$ is associated with a set of textual description $\mathcal{X}_{e}=\{X^{t}_{e},X^{d}_{e}\}$, where $X^{t}_{e}$ is its term and $X^{d}_{e}$ is its definition. Meanwhile, each directed edge $h=\langle p,c\rangle\in\mathcal{H}$ represents a parent-child relationship that points to a child entity $c$ from its most exact hypernymy entity $p$.
Definition 2 (Taxonomy Expansion)
Given a set of emerging conceptual entities $\mathcal{E}^{\prime}$, taxonomy expansion aims to incorporate these entities into an existing seed taxonomy $\mathcal{T}_{0}=(\mathcal{E}_{0},\mathcal{H}_{0})$. The goal is to expand $\mathcal{T}_{0}$ to be a larger taxonomy $\mathcal{T}=(\mathcal{E}_{0}\cup\mathcal{E}^{\prime},\mathcal{H}^{\prime})$. To insert each query entity $q\in\mathcal{E}^{\prime}$, we identify an appropriate anchor entity $a\in\mathcal{E}_{0}$, and introduce a new edge $\langle q,a\rangle$. Consequently, the updated edge set is $\mathcal{H}^{\prime}=\mathcal{H}_{0}{\cup}_{q\in\mathcal{E}^{\prime}}\{\langle q,a\rangle\}$.
3 Methodology
In this section, we provide a comprehensive overview of our proposed CodeTaxo designed for addressing the taxonomy expansion task. Specifically, CodeTaxo expands the existing taxonomy by prompting LLMs with code language. The pipeline of CodeTaxo is shown in Figure 2. Our CodeTaxo consists of three parts: Task Instruction, Taxonomy Code Representation, and Few-shot Demonstrations Construction.
3.1 Task Instruction
To enhance the effectiveness and accuracy of LLMs in completing the taxonomy expansion task, we propose a detailed task description along with a set of fundamental rules, denoted as $\mathcal{R}$, for expanding the existing taxonomy via the query entity. As illustrated in Figure 3, component (a) outlines the objectives of the taxonomy expansion task, framing it as a code completion task and specifying add_parent function should be employed. In component (b), we emphasize a set of fundamental rules $\mathcal{R}$ for the taxonomy expansion task. These rules include the following: 1. Do not use entities that are not covered in the existing taxonomy $\mathcal{T}_{0}=(\mathcal{E}_{0},\mathcal{H}_{0})$ ($\mathbf{r}_{1}$); 2. Maintain the output generation format by LLMs, consisting of one line of code followed by one line explaining why the model made that prediction ($\mathbf{r}_{2}$); 3. Refrain from generating additional content ($\mathbf{r}_{3}$). Additionally, the rule for generating an explanation for the prediction in $\mathbf{r}_{2}$ is optional for future analysis. In CodeTaxo, this rule is omitted as generating explanations is not required.
Figure 4: Python code in CodeTaxo defining a Entity class for managing parent-child relations.
3.2 Taxonomy Code Representation
To represent the existing taxonomy $\mathcal{T}_{0}=(\mathcal{V}_{0},\mathcal{E}_{0})$ as code language, we concatenate the entity class definition, representation of existing taxonomic relations, and the code completion prompt. We use Python as the programming language for the code prompt due to its widespread popularity.
3.2.1 Entity Class Definition
First, we define a base type Entity to be inherited by each entity mentioned in the taxonomy expansion. In Figure 4, we define a Python class named Entity that models a taxonomic structure with parent-child relations. The first line imports the List type from the typing module, which is used for type hinting. This allows the child attribute to be explicitly declared as a list of Entity objects.
The Entity class encapsulates the attributes and methods for managing hierarchical entities. The __init__ method initializes an instance of the Entity class with the following parameters:
  • name: A string storing the term of the entity.
  • description: A string storing the textual description of the entity
  • parent: An instance of the Entity class, denoting the parent entity within the taxonomy.
  • child: A list of entities, each an instance of the Entity class, storing the entity’s children.
These instance attributes are assigned as follows: self.name, self.description, self.parent, and self.child. Additionally, since we consider that each entity in the taxonomy should only have one parent entity, we do not use the List type for the parent attribute, unlike the child attribute.
The Entity class includes two methods for modifying the parent-child relations between entities. The first method, add_parent, assigns a parent entity to the current entity. It takes one parameter, parent, which is an instance of the Entity class. The second method, add_child, appends the child entity to the self.child list of the current entity. This method also requires one parameter, child, which is an instance of the Entity class.
3.2.2 Representing the Existing Taxonomy
To facilitate the taxonomy expansion, the initial taxonomy $\mathcal{T}_{0}$ is encoded using a programming language. Instances of the Entity class, as defined in Section 3.2.1, are created for each entity $e$ in the set $\mathcal{E}_{0}$ of $\mathcal{T}_{0}$. The taxonomy $\mathcal{T}_{0}$ is traversed from top to bottom, and for each entry, an entity $e\in\mathcal{E}_{0}$ is instantiated as follows:
$e=\operatorname{Entity}(\text{name}$ $={X}^{t}_{e},\text{description}={X}^{d}_{e},$
parent $=p_{e},\text{child}=\mathcal{C}_{e})$
where $p_{e}$ is the parent entity of $e$, and $\mathcal{C}_{e}=[c^{1}_{e},c^{2}_{e},\dots,c^{n}_{e}]$ is the list of its child entities.
Dataset SemEval-Sci SemEval-Env SemEval-Food WordNet Graphine
Metric Acc Wu&P Acc Wu&P Acc Wu&P Acc Wu&P Acc Wu&P
Self-supervised Setting
TaxoExpan 27.8 57.6 11.1 54.8 27.6 54.2 19.8 64.8 24.5 65.9
STEAM 36.5 68.2 36.1 69.6 34.2 67.0 23.2 62.4 20.3 63.1
HEF 53.6 75.6 55.3 71.4 47.9 73.5 16.4 60.3 25.5 66.5
Musubu 44.9 76.2 45.3 65.4 42.3 72.4 28.5 64.0 35.4 75.2
TEMP 57.8 85.3 49.2 77.7 47.6 81.0 29.4 65.7 35.9 73.8
BoxTaxo 31.8 64.7 38.1 75.4 31.4 66.8 26.4 63.9 29.2 68.2
TaxoPrompt 61.4 85.6 57.4 83.6 53.2 83.1 40.3 71.5 33.9 74.4
TaxoInstruct 45.9 76.2 48.8 77.2 34.3 70.2 43.3 71.8 31.8 69.0
1-shot Setting
NL (GPT-4o) 54.8 88.3 52.5 81.3 55.5 85.6 72.2 90.7 69.8 89.1
CodeTaxo (GPT-4o) 67.7 89.2 62.5 86.1 58.1 85.3 74.5 91.3 72.9 91.0
NL (GPT-4o-mini) 50.0 83.0 35.0 76.1 55.1 87.2 60.1 86.0 58.3 85.2
CodeTaxo (GPT-4o-mini) 58.1 85.6 42.5 76.0 55.9 85.3 68.8 89.2 61.5 85.1
5-shot Setting
NL (GPT-4o) 56.5 84.3 60.0 85.5 52.5 86.9 72.2 90.1 69.3 90.0
CodeTaxo (GPT-4o) 66.1 88.0 67.5 87.0 60.2 85.7 76.5 91.9 77.6 93.4
NL (GPT-4o-mini) 53.2 84.8 42.5 80.2 57.2 87.6 63.4 87.3 63.5 88.6
CodeTaxo (GPT-4o-mini) 59.7 84.8 47.5 78.3 58.9 87.9 66.8 88.6 70.3 89.1
Table 1: Performance on taxonomy expansion across two small-scale taxonomies (WordNet and Graphine) and three large-scale taxonomies (SemEval2016: science, environment, food). Bold indicates the highest score; underlined indicates the second-highest. All metrics are in percentages (%).
3.2.3 Semantic Similarity Filter
Including all entity $e\in\mathcal{E}_{0}$ to represent the existing taxonomy $\mathcal{T}_{0}$ presents two problems. First, large-scale taxonomies overload the LLM’s limited context window. Second, it unnecessarily expands the search space, introducing irrelevant entities and redundant information. To mitigate these issues, we propose a Semantic Similarity Filter that selects only entities relevant to the query $q$ for inclusion in the prompt context.
To compute the similarity between a query entity $q$ with its descriptive text $\mathcal{X}q=\{X^{t}_{q},X^{d}_{q}\}$ and an entity $e_{i}\in\mathcal{E}0$ with its descriptive text $\mathcal{X}{e_{i}}=\{X^{t}_{e_{i}},X^{d}_{e_{i}}\}$, we employ the pre-trained language model (PLM) as textual encoder. We concatenate the query entity $q$ and the $i$-th entity $e_{i}$ with special tokens [CLS] and [SEP], then encode the sequence using a pre-trained SimCSE model Gao et al. (2021). SimCSE converts them into $m$-dimensional representation $\mathbf{q},\mathbf{e}_{i}\in\mathbb{R}^{m}$:
[…]
The semantic relevance is calculated using cosine similarity between $\{\mathbf{e}_{i}\}^{n}_{i=1}$ and $\mathbf{q}$. We select the Top-$k$ entities most similar to query entity $q$ from the entity set $\mathcal{E}_{0}$ in $\mathcal{T}_{0}$ as follow:
$\mathcal{I}$ $=\underset{\begin{subarray}{c}\mathcal{I}\subseteq\{1,2,\ldots,n\},\\ |\mathcal{I}|=k\end{subarray}}{\operatorname{argmax}}\sum_{i\in\mathcal{I}}\text{cos\_sim}(\mathbf{e}_{i},\mathbf{q})$
where $\mathcal{I}$ is the index set of the selected entities $\mathcal{E}_{sel}=\{e_{i}|i\in\mathcal{I}\}$ that represents the existing taxonomy. $k$ is set to 50% of the entities in $\mathcal{E}_{0}$.
3.2.4 Code Completion Prompt.
The code completion prompt involves the instantiation of a query entity $q$ as an instance of the Entity class, as defined in Section 3.2.1. Since the query entity $q$ lacks information about its parent and child entities, it is instantiated as follows:
$q=\operatorname{Entity}(\text{name}$ $={X}^{t}_{q},\text{description}={X}^{d}_{q},$
[…]
Here, $X^{t}_{q}$ and $X^{d}_{q}$ define the query’s name and description, while None and [] indicate the absence of parent and child entities.
We include the requirement “Find the parent of the query node” as a comment to guide LLMs in selecting an anchor entity $a\in\mathcal{E}_{sel}$ as the parent entity for entity $q$. The output is the query $q$, an instance of the Entity class, which invokes the predefined method add_parent() to assign $a$ as its parent entity like q.add_parent(a).
We propose incorporating an optional feature in the code completion prompt: “then generating a comment to explain why it is the parent of the given query node”. This feature allows the LLM to simultaneously generate both the prediction and its rationale, improving explainability and revealing interesting insights, as discussed in Section 4.5.
3.3 Few-shot Demonstration Construction
To enhance LLMs’ ability to expand our existing taxonomy, we propose a method for constructing demonstrations using the initial taxonomy $\mathcal{T}_{0}$. Our demonstration selection strategy focuses on the semantic similarity between the query entity $q$ and entities $e\in\mathcal{E}_{0}$ in the existing taxonomy. Specifically, we use SimCSE encoding to calculate these similarities, selecting the top-5 entities from the existing set $\mathcal{E}_{0}$ based on their similarity to $q$:
$\mathcal{I}_{d}$ $=\underset{\begin{subarray}{c}\mathcal{I}_{d}\subseteq\{1,2,\ldots,n\},\\ |\mathcal{I}_{d}|=5\end{subarray}}{\operatorname{argmax}}\sum_{i\in\mathcal{I}_{d}}\text{cos\_sim}(\mathbf{e}_{i},\mathbf{q})$
Here, $\mathcal{I}_{d}$ represents the indices of entities selected for the demonstration set $\mathcal{E}_{demo}=\{e_{i}|i\in\mathcal{I}_{d}\}$. For each demonstration $d_{i}$, we treat each entity $e_{i}\in\mathcal{E}_{demo}$ as a query entity and, following the procedure outlined in Section 3.2.4, add its parent entity using the add_parent method.
4 Experiments
4.1 Experimental Settings
Datasets.
We evaluate taxonomy expansion on small-scale WordNet sub-taxonomies Bansal et al. (2014) and Graphine taxonomies Liu et al. (2021a). Additionally, we evaluate three large-scale taxonomies from SemEval-2016 Bordea et al. (2016) across science, environment, and food domains. For all benchmarks, 20% of leaf entities are reserved for testing, with the remaining entities used for training. See App. A.1 for details.
Baselines.
We evaluate CodeTaxo, using both GPT-4o and GPT-4o-mini, against self-supervised baselines including TaxoExpan Shen et al. (2020), STEAM Yu et al. (2020), HEF Wang et al. (2022), Musubu Takeoka et al. (2021), TEMP Liu et al. (2021b), BoxTaxo Jiang et al. (2023), TaxoPrompt Xu et al. (2022), and TaxoInstruct Shen et al. (2024), and prompting LLMs through natural language. See details in App. A.2.
Method Def. SemEval-Sci SemEval-Env SemEval-Food WordNet Graphine
Acc Wu&P Acc Wu&P Acc Wu&P Acc Wu&P Acc Wu&P
1-shot Setting
NL (GPT-4o) $\checkmark$ 54.8 88.3 52.5 81.3 55.5 85.6 72.2 90.7 69.8 89.1
$\times$ 59.7 89.0 57.5 82.8 56.4 87.0 68.1 89.1 68.8 90.1
CodeTaxo (GPT-4o) $\checkmark$ 67.7 89.2 62.5 86.1 58.1 85.3 74.5 91.3 72.9 91.0
$\times$ 56.5 84.5 55.0 85.1 56.8 86.1 66.4 88.4 69.8 88.8
5-shot Setting
NL (GPT-4o) $\checkmark$ 56.5 84.3 60.0 85.5 52.5 86.9 72.2 90.1 69.3 90.0
$\times$ 59.7 89.6 50.0 79.3 55.5 87.6 70.5 89.9 68.8 88.9
CodeTaxo (GPT-4o) $\checkmark$ 66.1 88.0 67.5 87.0 60.2 85.7 76.5 91.9 77.6 93.4
$\times$ 51.6 80.6 65.0 86.7 57.6 86.1 67.8 88.8 68.8 89.7
Table 2: Impact of Definition Sentences (Def.) on CodeTaxo and NL Performance in 1-Shot and 5-Shot Settings.
Evaluation Metrics:
We use two Accuracy (ACC) and Wu & Palmer similarity (Wu&P) to evaluate the performance of CodeTaxo and baselines. See details in App. A.3
Figure 5: Performance comparison of NL and CodeTaxo (CT) across Llama trained on Code and Natural Language domains. Due to limited contextual window sizes, evaluations were conducted on small-scale sub-taxonomies from WordNet (WN) and Graphine (G).
4.2 Experimental Results
4.2.1 Can CodeTaxo expand taxonomy better than other baselines?
We evaluate CodeTaxo against baseline methods for taxonomy expansion in Table 1, including self-supervised and in-context learning approaches. On WordNet and Graphine, both NL and CodeTaxo significantly outperform self-supervised baselines. In one-shot settings, CodeTaxo improves accuracy by 72.06% and 103.06% over the best self-supervised methods, demonstrating that minimal annotated data effectively unlocks LLMs’ internal knowledge, while self-supervised methods struggle with limited-scale taxonomies. On large-scale SemEval-2016 taxonomies, CodeTaxo surpasses the best self-supervised baseline, TaxoPrompt, by 10.26%, 8.89%, and 9.21% on SemEval-Sci, SemEval-Env, and SemEval-Food, respectively. While the NL prompt underperforms TaxoPrompt on SemEval-Sci and SemEval-Env, it exceeds TaxoPrompt on SemEval-Food but still trails CodeTaxo by 4.68%, highlighting CodeTaxo’s superior ability to capture taxonomic structures. Performance depends on LLM capability and demonstration count, with GPT-4o outperforming GPT-4o-mini and more demonstrations improving accuracy across benchmarks, underscoring the value of high-quality demonstrations for taxonomy expansion.
Setting Config. SemEval-Sci SemEval-Env SemEval-Food
Demo. Filter Acc Wu&P Acc Wu&P Acc Wu&P
1-shot $\times$ $\times$ 50.0 84.0 47.5 81.1 56.4 85.3
$\times$ $\checkmark$ 61.3 84.4 55.0 83.2 54.2 84.6
$\checkmark$ $\times$ 61.3 85.9 47.5 79.0 57.2 86.7
$\checkmark$ $\checkmark$ 67.7 89.2 62.5 86.1 58.1 85.3
5-shot $\times$ $\times$ 58.1 86.0 55.0 82.8 56.4 86.5
$\times$ $\checkmark$ 59.7 84.7 57.5 83.7 58.5 85.8
$\checkmark$ $\times$ 61.3 88.5 55.0 85.3 57.6 86.5
$\checkmark$ $\checkmark$ 66.1 88.0 67.5 87.02 60.2 85.7
Table 3: Ablation Study of two major modules in the CodeTaxo: All metrics are presented in percentages (%). Configurations indicate whether Demonstration Selection (Demo.) and Semantic Similarity Filter (Filter) were employed.
4.2.2 How does CodeTaxo perform across different large language models?
We evaluate CodeTaxo, a prompting method specifically designed for programming languages, by comparing its effectiveness against natural language prompting on both general-purpose LLMs and Code-LLMs. We used Llama-family models, including LLaMa-3-70B-instruct, CodeLLaMA-70B-instruct, and the smaller CodeLLaMA-34B-instruct, to evaluate how model size affects performance.  Given the limited contextual capacity of these models, we focused our evaluation on WordNet and Graphine, as shown in Figure 5. The results highlight CodeTaxo’s superior accuracy and Wu&P scores across all tested models, outperforming natural language prompts in representing taxonomic structures for black-box LLMs like GPT-4 and open-source LLMs. The analysis further reveals that CodeTaxo benefits more significantly from Code-LLMs, with a 13.33% accuracy improvement on WordNet compared to 6.51% for natural language prompts when transitioning to code language prompting. Notably, CodeLLaMA-34B-instruct, despite being smaller, showed better performance on WordNet and Graphine, emphasizing CodeTaxo’s efficiency and robustness.
Figure 6: Effect of Top-K relevant entities selected through SimCSE-based Semantic Similarity Filter.
4.3 Hyperparmeter Analysis of CodeTaxo
This section explores the impact of selecting Top-K entities using the Semantic Similarity Filter on model performance, with experiments conducted using GPT-4o-mini across three SemEval taxonomies. As shown in Figure 6, increasing the number of Top-K entities generally enhances performance by retaining more entities, thereby reducing the likelihood of filtering out the ground truth and boosting prediction accuracy. However, this improvement involves a trade-off: a smaller search space sharpens the model’s focus but increases the risk of excluding the ground truth. For instance, in the SemEval-Sci taxonomy, the model achieved optimal performance with a Hit@25 score of 78% by retaining the top 25 entities, demonstrating the filter’s ability to balance search space and coverage. To further refine this balance, we retained the top 50% of entities in our experiments, ensuring that Hit@n exceeded 90% across all benchmarks.
Query NL Output CodeTaxo Output CodeTaxo Exp.
functional genomics
Def. functional genomics is a field of molecular biology that attempts to describe gene functions and interactions
genetics
Def. genetics is a branch of biology concerned with the study of genes, genetic variation, and heredity in organisms.
genomics
Def. genomics is an interdisciplinary field of biology focusing on the structure, function, evolution, mapping, and editing of genomes
genomics
Exp. Functional genomics is a subfield of genomics that focuses on understanding the functions and inter- actions of genes within the genome.
geostrategy
Def. geostrategy, a subfield of geopolitics, is a type of foreign policy guided principally by geographical factors as they inform, constrain, or affect political and military planning
politics
Def. politics is the set of activities that are associated with making decisions in groups, or other forms of power relations between individuals, such as the distribution of resources or status
geopolitics
Def. geopolitics on politics and international relations
geopolitics
Exp. : Geostrategy is a subfield of geopolitics, which focuses on geographic factors influencing political and military planning.
Table 4: Case study comparing the outputs of the CodeTaxo and NL prompt using the SemEval-Sci benchmarks. The table presents the definitions (Def.) of each model’s prediction, additionally with the CodeTaxo explanations (Exp.) provided in the last column (CodeTaxo Exp.). Yellow highlights emphasize the specific focus of the query within its definition, as correctly captured by CodeTaxo, while green highlights indicate broader, less precise concepts used by the NL model.
4.4 Ablation Study
4.4.1 Insight of Definition Sentences
We performed an ablation study on definition sentences, a vital data source for taxonomy expansion tasks, using two prompting methods: NL and CodeTaxo. Our results in Table 2 show that without definition sentences, CodeTaxo suffers a substantial drop in accuracy and Wu&P across all benchmarks in both 1-shot and 5-shot settings, highlighting its reliance on semantic information from definitions to establish taxonomic relationships. Interestingly, NL performed better without definition sentences in specific benchmarks (SemEval-Sci, SemEval-Env, SemEval-Food) in the 1-shot setting, and in SemEval-Sci and SemEval-Food in the 5-shot setting. This suggests that NL struggles to process definition information effectively, potentially leading to incorrect predictions when overloaded with definitional content.
4.4.2 Effevtiveness of Demo. Selection and Semantic Similarity Filter
We performed an ablation study on the three SemEval2016 benchmarks mentioned above to assess the effectiveness of the two primary modules in CodeTaxo: Demonstration Selection (Demo.) and the Semantic Similarity Filter (Filter). Due to the relatively small size of the taxonomies in WordNet and Graphine, filtering redundant entities from the existing taxonomies was unnecessary. The results, presented in Table 3, indicate that selecting demonstrations related to the query entity and filtering out unrelated entities in the existing taxonomy significantly improves taxonomy expansion. This finding suggests that incorporating more relevant contextual information and reducing redundant information to narrow the search space is beneficial for both accuracy and the Wu&P score across all SemEval2016 benchmarks.
4.5 Case Study
This section presents a case study demonstrating the effectiveness of our CodeTaxo framework by comparing its outputs to those of the natural language (NL) prompt, alongside model predictions and corresponding definitions in Table 4. Notably, CodeTaxo aligns closely with the ground truth and generates explanations using the prompt from Section 3.2.4 to facilitate insightful discussions. For instance, in the query functional genomics, CodeTaxo accurately classifies it under genomics, emphasizing its focus on the “functions and interactions of genes within the genome”, whereas the NL model incorrectly selects the broader term genetics. Similarly, CodeTaxo identifies geopolitics as the parent entity of geostrategy, highlighting its emphasis on geographic factors, while the NL model selects the more general category of politics. These cases showcase CodeTaxo’s ability to leverage definition information for a comprehensive understanding of taxonomy structures, resulting in more precise predictions.
5 Related Works
5.1 Taxonomy Expansion
In taxonomy expansion, various approaches have been developed to integrate emerging entities into existing taxonomies. Aly et al. Aly et al. (2019) and Ma et al. utilized hyperbolic embeddings to capture taxonomic relations, while Jiang et al. Jiang et al. (2023) and Xu et al. Xu et al. (2024) employed box embedding and fuse embedding instead of single vector embedding to encode taxonomic relations respectively. Manzoor et al. Manzoor et al. (2020) introduced implicit edge semantics to enhance entity representations. Self-supervised methods, such as Egonet Shen et al. (2020), mini-path Yu et al. (2020), and Ego-Tree Wang et al. (2021), have also been explored to model structural information within taxonomies. To leverage more semantic information from the textural description of entities, Liu et al. Liu et al. (2021b), Takeoka et al. Takeoka et al. (2021) and Xu et al. Xu et al. (2022) fine-tuned BERT-based models to leverage textual descriptions of entities. Zhu et al. Zhu et al. (2023) integrates textual and visual semantics to capture the hierarchical relation between entities. Shen et al. Shen et al. (2024) and Moskvoretskii et al. Moskvoretskii et al. (2024) unified framework combining various taxonomy construction tasks for instruction tuning. To our knowledge, CodeTaxo is the first work to perform taxonomy expansion via prompting LLMs.
5.2 Code-LLMs for Structured Tasks
Recent studies have demonstrated the strong performance of Code-LLMs in complex reasoning tasks Yang et al. ; MA et al. , including symbolic reasoning Madaan et al. (2022); Cheng et al. , graph reasoning Cai et al. (2024), event structure prediction Wang et al. (2023); Chen et al. (2023), mathematical reasoning Gao et al. (2023), and knowledge graph construction Li et al. (2023); Bi et al. (2024). These works highlight Code-LLMs’ ability to transform unstructured text into structured representations, enabling advanced reasoning tasks. In this paper, we focus on enhancing Code-LLMs’ ability to comprehend and expand existing taxonomies through emerging query entities.
6 Conclusion
In this paper, we introduce CodeTaxo, a novel approach to taxonomy expansion that leverages code-based prompts to effectively utilize the inherent knowledge within LLMs. Our method addresses key challenges in taxonomy expansion by reformulating the task as a code completion problem and employing a Semantic Similarity Filtering mechanism to optimize the use of LLMs’ contextual capacity. Extensive experiments on small-scale and large-scale taxonomies demonstrate that CodeTaxo achieves state-of-the-art performance in both one-shot settings and five-shot settings. We envision CodeTaxo as a powerful framework for integrating emerging entities into existing taxonomies by accurately identifying appropriate parent entities and also providing new insights for leveraging LLMs in structured knowledge tasks.
Limitations
This study represents an initial effort to utilize LLMs for taxonomy expansion. Our primary objective is to identify an effective in-context learning strategy to leverage the potential of LLMs. We acknowledge that the performance and scalability of CodeTaxo are constrained by the inherent knowledge of LLMs and the limitations of their context window size. While this paper does not address the challenges of expanding LLM knowledge or increasing context window size, we hope that our work will inspire further research in these areas.
Ethics Statement
Our research addresses taxonomy expansion within general knowledge domains, leveraging our proposed method, CodeTaxo, which uses large language models (LLMs) to generate structured knowledge and overcome the limitations of traditional manual taxonomy construction. We exclusively utilize publicly available datasets and benchmarks, avoiding user-generated, private, or sensitive data to ensure compliance with privacy and ethical standards. While our datasets do not engage directly with ethically sensitive content, LLMs inherently carry biases from their pre-training data, which may influence the structure and content of the expanded taxonomies. To address this, we integrate mechanisms for generating explanatory outputs, enabling detailed scrutiny of the model’s reasoning and identifying potential biases. Additionally, we recognize the risks of applying similar methodologies to subjective or sensitive domains, which could lead to misrepresentation or bias. To mitigate such risks, we emphasize collaboration with domain experts and advocate for responsible application of our methodologies across diverse fields, aiming to promote fairness, accuracy, and ethical research practices.
Appendix A Appendix
A.1 Datasets
We evaluate the performance of taxonomy expansion methods on small-scale taxonomies using WordNet Sub-taxonomies from Bansal et al. (2014), and Graphine taxonomies from Liu et al. (2021a). Specifically, we use 35 Graphine taxonomies with fewer than 100 entities, selected from a total of 227 taxonomies. For the Graphine dataset, we selected 35 taxonomies with fewer than 100 entities out of 227 total taxonomies. In our experiment with WordNet, we utilized 114 sub-taxonomies from the test sets. Additionally, we evaluate three large-scale taxonomies from SemEval-2016 Bordea et al. (2016) across science, environment, and food domains. Table 5 presents the statistics of these taxonomies, all of which contain entities and definitions curated by human experts. For all benchmarks, 20% of leaf entities are reserved for testing, with the remaining entities used for training.
#Concepts #Edges Depth License WordNet 20.5 19.5 3.0 WordNet Graphine 48.2 48.2 4.6 None SemEval-Sci 429.0 451.0 8.0 None SemEval-Env 261.0 261.0 6.0 None SemEval-Food 1,486.0 1,576.0 8.0 None
Table 5: Statistics of five taxonomy benchmarks. For WordNet and Graphine, we report the average for taxonomies included in these two benchmarks.
A.2 Baselines
We compare our method with the following baselines for taxonomy expansion, all experiments are implemented in a server with three NVIDIA A6000 GPUs:
  • TaxoExpan Shen et al. (2020): adopts GNNs to encode local ego-graphs in taxonomy to enhance entity representation.
  • STEAM Yu et al. (2020): utilizes the mini-path information to capture the global structure of the taxonomy.
  • HEF Wang et al. (2022): represents taxonomies as ego-trees to capture hierarchy, fully leveraging the hierarchical structure to improve taxonomy coherence.
  • Musubu Takeoka et al. (2021): leverages pre-trained models and fine-tunes them as sentence classifiers using queries generated from Hearst patterns.
  • TEMP Liu et al. (2021b): utilizes a pre-trained model to encode text descriptions of each concept in the taxonomy. It incorporates taxonomic structure information through taxonomy paths.
  • BoxTaxo Jiang et al. (2023): represent the entities via box embeddings instead of single vector embeddings to capture the hierarchical relation between entities.
  • TaxoPrompt Xu et al. (2022): adopt prompt tuning on the BERT-based encoder model to capture the taxonomic structure.
  • TaxoInstruct Shen et al. (2024): a unified framework for taxonomy-related tasks using instruction tuning, focused solely on taxonomy expansion for fair comparison.
To the best of our knowledge, CodeTaxo represents the first work to address taxonomy expansion using an in-context learning approach. To validate the effectiveness of the code language based prompt design, we additionally propose a prompting method based on natural language prompts. The results obtained using the natural language prompt (NL) are presented in Table 1. To ensure that the natural language prompt communicates the same information as code language prompt in CodeTaxo, we represent each entity using natural language to describe its surface name, definition, parent, and children list. The details of the NL prompt are provided in Table 7. For a more direct comparison, we also demonstrate CodeTaxo’s predictions on the same example in Table 8.
A.3 Evaluation Metrics.
The performance of CodeTaxo and the baseline models for taxonomy expansion tasks is evaluated using commonly adopted metrics, including accuracy (Acc) and Wu & Palmer similarity (Wu&P), as established in prior work Yu et al. (2020); Liu et al. (2021b); Wang et al. (2021). Since CodeTaxo is a generation-based method rather than a ranking-based one, the mean reciprocal rank (MRR) used in the baselines is not applicable to CodeTaxo.
Dataset 1-shot 5-shot NL CodeTaxo NL CodeTaxo SemEval-Sci 15737.2 9701.4 16095.1 10342.6 SemEval-Env 8965.7 5693.6 9325.0 6321.1 SemEval-Food 48908.1 30536.5 49266.4 31176.3 WordNet 948.9 1369.2 1306.4 1962.3 Graphine 2486.0 3223.9 2855.2 3893.3
Table 6: Comparison of average tokens used by NL and CodeTaxo across 5 benchmarks in 1-shot and 5-shot settings.
A.4 Efficiency Analysis of CodeTaxo
Token Consumption
Table 6 compares average token usage across benchmarks and prompt types (CodeTaxo vs. NL) in 1-shot and 5-shot settings. The findings highlight CodeTaxo’s efficiency in reducing token usage while maintaining effectiveness. Notably, in SemEval2016, CodeTaxo cuts token usage by approximately 37.6% in the SemEval-Food task compared to natural language prompts. However, in the WordNet and Graphine datasets, CodeTaxo uses slightly more tokens due to the need to define Entity classes and methods. Overall, the significant reduction in token usage in SemEval2016 underscores CodeTaxo’s efficiency, especially in contexts with limited token windows.
Natural Language Prompt
User: Given the current taxonomy, find the parent of the query node. Please note that the query node may be a new node not in the current taxonomy. The parent of given query node always exists, so do not generate ’none’ or ’not found’. You only need to answer the entity name and do not generate any additional content or comments.
lunacy: obsolete terms for legal insanity; parent: insanity; children: [].
irrationality: the state of being irrational; lacking powers of understanding; parent: insanity; children: [].
dementia: mental deterioration of organic or functional origin; parent: insanity; children: [’presenile dementia’, ’alcoholic dementia’, ’senile dementia’].
alcoholic dementia: dementia observed during the last stages of severe chronic alcoholism; involves loss of memory for recent events although long term memory is intact; parent: dementia; children: [].
Pick’s disease: a progressive form of presenile dementia found most often in middle-aged and elderly women and characterized by degeneration of the frontal and temporal lobes with loss of intellectual ability and transitory aphasia; parent: presenile dementia; children: [].
derangement: a state of mental disturbance and disorientation; parent: insanity; children: [].
craziness: informal terms for insanity; parent: insanity; children: [].
presenile dementia: dementia with onset before the age of 65; parent: dementia; children: ["Pick’s disease"].
senile dementia: dementia of the aged; results from degeneration of the brain in the absence of cerebrovascular disease; parent: dementia; children: [].
insanity: relatively permanent disorder of the mind; parent: None; children: [’irrationality’, ’dementia’, ’craziness’, ’derangement’, ’lunacy’].
Query node: Alzheimer’s disease
The parent of query node:
Assistant: dementia
Ground Truth: presenile dementia
Table 7: Example of Natural Language (NL) Prompt.
Code Prompt
User: Complete the next line of code according to the comments and the given code snippet. You need to find the parent of the query node in the given current taxonomy and use the add_parent function. The parent of given query node always exists in the given current taxonomy, so do NOT generate node that is NOT in the given current taxonomy. Note that you only need to complete the next ONE line of code, do not generate any additional content or comments.
from typing import List
class Entity: def __init__(self, name: str, description: str, parent: str, child: List[’Entity’]): self.name = name self.description = description self.parent = parent self.child = child def add_parent(self, parent: ’Entity’): self.parent = parent.name parent.add_child(self) def add_child(self, child: ’Entity’): self.child.append(child)
# Creating entities and establishing parent-child relationship lunacy = Entity(name=’lunacy’, description=’obsolete terms for legal insanity’, parent=insanity, child=[]) irrationality = Entity(name=’irrationality’, description=’the state of being irrational; lacking powers of understanding’, parent=insanity, child=[]) dementia = Entity(name=’dementia’, description=’mental deterioration of organic or functional origin’, parent=insanity, child=[’presenile dementia’, ’alcoholic dementia’, ’senile dementia’]) alcoholic_dementia = Entity(name=’alcoholic dementia’, description=’dementia observed during the last stages of severe chronic alcoholism; involves loss of memory for recent events although long term memory is intact’, parent=dementia, child=[]) Pick’s_disease = Entity(name=’Pick’s disease’, description=’a progressive form of presenile dementia found most often in middle-aged and elderly women and characterized by degeneration of the frontal and temporal lobes with loss of intellectual ability and transitory aphasia’, parent=presenile dementia, child=[]) derangement = Entity(name=’derangement’, description=’a state of mental disturbance and disorientation’, parent=insanity, child=[]) craziness = Entity(name=’craziness’, description=’informal terms for insanity’, parent=insanity, child=[]) presenile_dementia = Entity(name=’presenile dementia’, description=’dementia with onset before the age of 65’, parent=dementia, child=["Pick’s disease"]) senile_dementia = Entity(name=’senile dementia’, description=’dementia of the aged; results from degeneration of the brain in the absence of cerebrovascular disease’, parent=dementia, child=[]) insanity = Entity(name=’insanity’, description=’relatively permanent disorder of the mind’, parent=None, child=[’irrationality’, ’dementia’, ’craziness’, ’derangement’, ’lunacy’])
# creating query node Alzheimer’s_disease = Entity(name=’Alzheimer’s disease’, description=’a progressive form of presenile dementia that is similar to senile dementia except that it usually starts in the 40s or 50s; first symptoms are impaired memory which is followed by impaired thought and speech and finally complete helplessness’, parent=None, child=[])
# Finding the parent of query node
Assistant: Alzheimer’s_disease.add_parent(presenile_dementia)
Ground Truth: presenile dementia
Table 8: Example of Code-based Prompt.